What is the equation of an ellipse with vertices (+-2, 0) and (0,+-3)
Answer by Boreal(15235)

The major axis is the longer and the square of a, its length, goes under the appropriate variable. Here, the major axis is along the y-axis from -3 to +3.
a=3 and a^2=9. This will be a vertical ellipse on its long axis and centered at the origin.
The minor axis is along the x-axis from -2 to 2, so b is 2 and b^2=4
x^2/4+y^2/9=1
or 9x^2+4y^2=36